Preview of Homework 1 R.1 The probability distribution for diff R.2 The probability distribution for max R.3 Expected value of diff R.4 Expected value of max R.6 Expected diff squared R.7 Expected max squared R.9 Variance of diff R.10 Variance of max Histogram for max Histogram for diff4. Now we return to the red and green dice from Homework 1. Use the lines below to simulate the random variables from Doughertys Chapter 1 problems, and then use R to estimate the answers to each of the problems below. ## Generating the Random Variables for Pages 13 and 14 red - sample (c(1:6),repl-T, size-10000) green = sample(c(1:6),repl-T,size-10000) sum = red-green max - ifelse (red > green, red, green) diff abs(red-green) table (sum) ## sum ## 2 3 4 5 6 7 8 9 10 11 12 ## 281 605 806 1075 1333 1679 1401 1123 841 576 280 mean (sum) ## [1] 7.0148 mean (sum 2) ## [1] 55.1342 var (sum) ## [1] 5.927374 sd (sum) ## [1] 2.43462 sum( (sum-mean (sum)) 2/9999 ## [1] 5.927374 mean (sum 2)-mean (sum)2 ## [1] 5.926781 Why doesnt the last expression equal the other variances? (Because mean divides by n and not n - 1 hist (sum)Histogram of sum 4 10 12 sum

Need help with the codings in R Markdown (in R Studio)
Answer #1

Sol: Since sample mean and sample variance is unbiased estimator of population mean and population variance respectively.

E(ar{x})=mu    and

E(s^2)=sigma ^2

Where,

ar{x} is a sample mean and mu is a population mean.

s^2=rac{1}{n-1}sum (x_{i}-ar{x})^2 is sample variance.

(ri -T) is population variance.

Therefore, R-code result gives approximately same value of sample variance (dividing by n-1) and population variance (dividing by n)

> red=sample(c(1:6),repl=T,size=10000)
> green=sample(c(1:6),repl=T,size=10000)
> sum=red+green
> max=ifelse(red>green, red, green)
> diff=abs(red-green)
> table(sum)
sum
   2    3    4    5    6    7    8    9   10   11   12
270 555 850 1132 1380 1649 1372 1084 848 591 269
> mean(sum)
[1] 7.0029
> mean(sum^2)
[1] 54.9115
> var(sum)
[1] 5.871479
> sd(sum)
[1] 2.423113
> sum((sum-mean(sum))^2)/9999 # This case provides sample variance.
[1] 5.871479
> mean(sum^2)-mean(sum)^2    # This case provides population variance.
[1] 5.870892
> hist(sum)